Adipic acid is an inorganic compound composed of 49.31% C, 43.79% O, and the test hydrogen. If the molar mass of adipic acid is 146.1 g/mol, what are the empirical and molecular formulas of adipic acid?

(I’ve already got most of the work done. I just don’t know how to get the empirical mass.)
1
Expert's answer
2019-08-29T03:34:53-0400

Let 100g of adipic acid is taken

Amount of C"=" "49.31g"

Amount of O"=" "43.79g"


Amount of H "=100-49.31-43.79=6.9g"


Moles of C"=\frac{49.31}{12.0079}=4.105 mole" C


Moles of H ="\frac{6.9}{1.0079}=6.845mole" H


Moles of O"=" "\frac{43.79}{15.999}=2.737mole" O

Find small, whole number ratio; divide all moles by 2.737

C: "\frac{4.105}{2.737}=" 1.5 moles C


H: "\frac{6.845}{2.737}=" 2.5 moles H


O: "\frac{2.737}{2.737}=" 1 mole O

Multiply by 2 in all moles for getting whole numbers

We get,

3 moles C

5 moles H

2 moles O

Then empirical formula becomes,

"C_3H_5O_2"

Mass of empirical formula of adipic acid "=73" "\frac{g}{mole}"

 the molar mass of adipic acid "=" 146.1 "\frac{g}{mole}"

Molar mass is two times the empiricl formula

So,

Molecular formula of adipic acid is "C_6H_{10}O_4"